OpenStudy (anonymous):

Verify (1-cot^2(w)+cos^2(w)*cot^2(w))/csc^2(w) = sin^4(w)

OpenStudy (anonymous):

\[(1-\cot ^{2}(w) + \cos ^{2}(w)* \cot^2(w)) / \csc^2(w) = \sin^4(w)\]

OpenStudy (anonymous):

[1-(cos^2w/sin^2w)+(cos^4w/sin^2w)]sin^2w=sin^2w-cos^2w+cos^4w

OpenStudy (anonymous):

I got \[\sin^2(w) - \cos^2(w) + \cos^4(w) = \sin^4(w)\] but not sure where to go from here

OpenStudy (anonymous):

does \[\sin^2(w) - \cos^2(w) = -1??\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

are you sure the identity is correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my homework problem in the book

OpenStudy (anonymous):

my answer book looks like this: \[(1-\cot^2(w)+\cos^2(w)\cot^2(w))/\csc^2(w)\] down to: \[(1-\cot^2(w)(1-\cos^2(w))/\csc^2(w)\] but i'm not sure how

OpenStudy (anonymous):

then \[(1-\cot^2(w)\sin^2(w))/\csc^2\] then: \[(1-(\cos^2(w)/\sin^2(w))\sin^2(w))/\csc^2(w)\] then \[(1-\cos^2(w))/\csc^2(w)\] then \[\sin^2(w)/\csc^2(w)\] to finally get \[\sin^4(w)\]